OpenStudy (anonymous):

8(x + 1) - 3(x + 4) = 7(2 - x)

OpenStudy (lilg132):

8x + 8 - 3x - 12 = 14 - 7x

OpenStudy (lilg132):

8x - 3x + 7x = 14 - 8 + 12 12x = 18 12x / 12 = 18 / 12 x = 1.5

OpenStudy (anonymous):

x = 1.5

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

1.5